The sun is shining down with an angle of depression of 75°. How long is the shadow of a boy who is 1.2 meters tall


1
Expert's answer
2021-05-10T17:17:11-0400

Angle of depression = Angle of elevation = 75°(Alternate angles are equal)


Angle of elevation = 75°

Height of the boy = 1.2m

Length of the boy's shadow = x m


"Tan∅=\frac{opposite}{adjacent}"

"Tan75°=\frac{1.2}{x}"

=> "x = \frac{1.2}{tan75°} = 0.3215"


Hence, the boy's shadow is roughly 0.32 metres long